# On the Lipschitz Continuity of Set Aggregation Functions and Neural Networks for Sets

### v3

@@ 1 Introduction
Prior work estimates mainly the Lipschitz constant of architectures composed of fully-connected and convolutional layers. However, in several domains, input data might correspond to complex objects which consist of other simpler objects. We typically model these complex objects as sets or multisets (i. e., a generalization of a set). For instance, in computer vision, a point cloud is a set of data points in the $3$-dimensional space. Likewise, in natural language processing, documents may be represented by multisets of word embeddings. Neural networks for sets typically consist of a series of fully-connected layers followed by an aggregation function which produces a representation for the entire multiset (Zaheer et al., 2017; Qi et al., 2017a). For the model to be invariant to permutations of the multiset’s elements, a permutation invariant aggregation function needs to be employed. Common functions include the sum, mean and max operators. While previous work has studied the expressive power of neural networks that employ the aforementioned aggregation functions (Wagstaff et al., 2019; 2022), their Lipschitz continuity and stability to perturbations remains underexplored. Besides those standard aggregation functions, other methods for embedding multisets have been proposed recently such as the Fourier Sliced-Wasserstein embedding which is bi-Lipschitz with respect to the Wasserstein distance (Amir and Dym, 2025).
@@ 3.1 Lipschitz Continuity of Aggregation Functions
The above theoretical result suggests that there is some correspondence between the three aggregation functions and the three distance functions for unordered multisets. In fact, each aggregation function seems to be closely related to a single metric. We also observe that while the Lipschitz constants of the sum and mean functions with respect to the matching distance and to EMD, respectively, is constant (equal to $1$), the Lipschitz constant of the max function with respect to the Hausdorff distance depends on the dimension $d$ of the vectors (which is typically larger than $1$). A direct consequence of the above Theorem is that no two of the considered distance functions are bi-Lipschitz equivalent.
@@ 4 Numerical Experiments
We experiment with two datasets from different domains: (i) ModelNet40: it contains 12,31112{,}311 3D CAD models that belong to $40$ object categories (Wu et al., 2015); and (ii) Polarity: it contains 10,66210{,}662 positive and negative labeled movie review snippets from Rotten Tomatoes (Pang and Lee, 2004). Note that the samples of both datasets can be thought of as multisets of vectors. Each sample of ModelNet40 is a multiset of $3$-dimensional vectors. Polarity consists of textual documents, and each document is represented as a multiset of word vectors. The word vectors are obtained from a publicly available pre-trained model (Mikolov et al., 2013).
@@ 4.1 Lipschitz Constant of Aggregation Functions
In the first set of experiments, we empirically validate the results of Theorem 3.1 and Lemma 3.2. To obtain a collection of multisets of vectors, we train three different neural network models on the ModelNet40 and Polarity datasets. The difference between the three models lies in the employed aggregation function: mean, sum or max. More details about the different layers of those models are given in Appendix D. Note that the multisets used to verify the Lipschitz constants of the aggregation functions could, in principle, be generated by any means. We use those models to create multisets since the objective is to investigate how these functions behave in comparison to the derived bounds when the inputs are sampled from real distributions, rather than artificially generated data that do not occur in practice. Once the models are trained, we feed the test samples to them. For each test sample, we store the multiset of vectors produced by the layer of the model that precedes the aggregation function, and we also store the output of the aggregation function (a vector for each multiset). We then randomly choose $100$ test samples, and for each pair of those samples, we compute the EMD, Hausdorff distance and matching distance of their multisets of vectors, and also the Euclidean distance of their vector representations produced by the aggregation function. This gives rise to $9$ combinations of distance functions and aggregation functions in total. Due to limited available space, we only show results for ModelNet40 in Figure 1. The results for Polarity can be found in Appendix E.1. Note that there are $\binom{100}{2}=4{,}950$ distinct pairs in total. Therefore, 4,9504{,}950 dots are visualized in each subfigure. To quantify the relationship between the output of the distance functions for multisets and the Euclidean distances of their vector representations, we compute and report the Pearson correlation coefficient. We observe from Figure 1 that the Lipschitz bounds (dash lines) successfully upper bound the Euclidean distance of the outputs of the aggregation functions. Note that all point clouds contained in the ModelNet40 dataset have equal cardinalities. Therefore, the conclusions of both Theorem 3.1 and Lemma 3.2 apply to this case, and thus we can derive Lipschitz constants for $7$ out of the $9$ combinations of distance functions for multisets and aggregation functions. We can see that the bounds that are associated with the mean and sum functions are tight, while those associated with the max function are relatively loose. We also observe that the distances of the representations produced by the mean and sum functions are very correlated with the distances produced by all three considered distance functions for multisets. On the other hand, the max function gives rise to representations that are less correlated with the produced distances.
@@ 4.2 Upper Bounds of Lipschitz Constants of Neural Networks for Sets
In the second set of experiments, we empirically validate the results of Theorem 3.4 and Lemma 3.5 on the ModelNet40 and Polarity datasets. We build neural networks that consist of three layers: (i) a fully-connected layer; (ii) an aggregation function; and (iii) a second fully-connected layer. Therefore, those models first transform the elements of the input multisets using an affine function, then aggregate the representations of the elements of each multiset and finally they transform the aggregated representations using a another affine function. Note that the Lipschitz constant of an affine function is equal to the largest singular value of the associated weight matrix, and can be exactly computed in polynomial time. We thus denote by $\text{Lip}(\text{FC}_{1})$ and $\text{Lip}(\text{FC}_{2})$ the Lipschitz constants of the two fully-connected layers, respectively. Note also that the Lipschitz constant of most activation functions (e. g., ReLU, LeakyReLU, Tanh) is equal to $1$. Therefore, we can compute an upper bound of the Lipschitz constant of some models using Theorem 3.4 and Lemma 3.5. To train the models, we add a final layer to them which transforms the vector representations of the multisets into class probabilities. We use the same experimental protocol as in subsection 4.1 above (i. e., we randomly choose $100$ test samples). We only provide results for ModelNet40 in Figure 2, while the results for Polarity can be found in Appendix E.2. Since all point clouds contained in the ModelNet40 dataset have equal cardinalities, the conclusions of both Theorem 3.4 and Lemma 3.5 apply to this setting, and thus, once again, we can derive upper bounds on the Lipschitz constants for $7$ out of the $9$ combinations of distance functions for multisets and aggregation functions. We observe that the dash lines (Lipschitz upper bounds from Theorem 3.4 and Lemma 3.5) indeed upper bound the Euclidean distance of the outputs of the aggregation functions. We can also see that the bounds that are associated with the mean function are tight, while those associated with the sum and max functions are relatively loose and very loose, respectively. We also observe that the distances of the representations produced by the mean and sum functions are very correlated with the distances produced by EMD and matching distance. On the other hand, the max function gives rise to representations whose distances are less correlated with the distances produced by the distance functions for multisets.
@@ Appendix A Related Work
The seminal work of Zaheer et al. (2017) introduced DeepSets, a model that uses the sum aggregator to produce permutation-invariant set representations. They showed that when the elements of the input sets come from a countable set $\mathcal{X}$, then for an appropriate $f\colon\mathcal{X}\rightarrow\mathbb{R}$, the function defined as $g(\{x_{1},\ldots,x_{n}\})=\sum_{i=1}^{n}f(x_{i})$ maps the input sets injectively to $\mathbb{R}$. This result was later extended to multisets (Xu et al., 2019). In the countable case, an embedding dimension of $1$ already suffices for injectivity. When $\mathcal{X}=\mathbb{R}$, multiset cardinalities are bounded by $m$ and $f$ is continuous, an embedding dimension of at least $m$ is both necessary and sufficient for injectivity (Wagstaff et al., 2019; 2022). For $\mathcal{X}=\mathbb{R}^{d}$, an embedding dimension of at least $md$ is necessary (Joshi et al., 2023). Most of these results rely on polynomial constructions to build injective multiset functions, after which the universal approximation theorem is invoked to argue that MLPs can approximate such polynomials. Amir et al. (2023) investigate whether MLP-based multiset functions are actually injective. They show that injectivity depends on the activation function. Specifically, analytic non-polynomial activation functions yield injective models, while networks with piecewise linear activation functions are injective only when $\mathcal{X}$ is finite or corresponds to certain irregular, countably infinite sets.
@@
Besides DeepSets, several other architectures and aggregation functions have been proposed recently. PointNet is another important architecture, primarily designed for point cloud data (Qi et al., 2017a). It consists of the same components as DeepSets, but instead of a sum aggregation function, it employs a max aggregator. To allow PointNet to capture local structures at different scales, Qi et al. (2017b) proposed PointNet++, a hierarchical model which applies PointNet recursively to nested partitions of the input set. Janossy pooling applies a neural network to all permutations of the input data and averages their outputs (Murphy et al., 2019). Since computing all permutations is generally intractable, the authors also propose some practical approximations. Set Transformer is a variant of the Transformer architecture designed for sets (Lee et al., 2019). Due to its attention mechanism, the model can capture interactions between elements in the input set. RepSet is another model designed for set-structured data which generates set representations by comparing the input set against learnable latent sets using a network flow algorithm (Skianis et al., 2020). FSPool sorts each feature across the elements of the set, and then computes a weighted sum of the elements where different weights can be learned for each feature dimension (Zhang et al., 2020). Pellegrini et al. (2021) propose a a learnable aggregation function which can approximate common aggregation functions (e. g., mean, sum, max), but also more complex functions. Kimura et al. (2024) introduce the Hölder’s Power DeepSets, a model tha generalized DeepSets by employing function known as power‑mean (a.k.a. Hölder mean), controlled by an exponent $p$.
@@ Appendix B Proofs
Thus, we only need to prove that the triangle inequality holds. Let $Z\in\mathcal{S}(\mathbb{R}^{d}\setminus\{\mathbf{0}\})$. The three multisets can have different cardinalities. Let $|X|=m$, $|Y|=n$ and $|Z|=k$. Then, $X=\{\!\!\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{m}\}\!\!\}$, $Y=\{\!\!\{\mathbf{u}_{1},\mathbf{u}_{2},\ldots,\mathbf{u}_{n}\}\!\!\}$ and $Z=\{\!\!\{\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{k}\}\!\!\}$. There are $6$ different cases. But it suffices to show that the triangle inequality holds when $|X|\geq|Y|\geq|Z|$, when $|Z|\geq|Y|\geq|X|$ and when $|X|\geq|Z|\geq|Y|$. Proofs for the rest of the cases are similar.
@@
Note that $d_{M}(X,Z)+d_{M}(Z,Y)$ can be written as a sum of $k$ terms, where each term corresponds to one of the above $4$ sums of norms. If we take pairs of terms of types $2$ and $3$ and we sum them, we have that:
@@
Note also that type $2$ occurs mnm-n times more than type $3$. Therefore, using the inequalities for the $4$ types of sums of norms above, we have:
@@ B.3.1 The mean function is Lipschitz continuous with respect to EMD
The mean function is thus Lipschitz continuous with respect to EMD and the Lipschitz constant is equal to $1$.
@@ B.3.2 The mean function is not Lipschitz continuous with respect to the matching distance
Suppose that the mean function is Lipschitz continuous with respect to the matching distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>(2L+1)\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1}\}$ be two multisets, consisting of $2$ and $1$ vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$, and $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$. Clearly, we have that $d_{M}(X,Y)=\|\mathbf{v}_{2}\|=\sqrt{d}\epsilon$. We also have that:
@@ B.3.3 The mean function is not Lipschitz continuous with respect to the Hausdorff distance
Suppose that the mean function is Lipschitz continuous with respect to the Hausdorff distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>3L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ be two multisets, consisting of $3$ and $2$ vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(-c,-c,\ldots,-c)^{\top}$, $\mathbf{v}_{2}=\mathbf{u}_{2}=(c,c,\ldots,c)^{\top}$ and $\mathbf{v}_{3}=(c+\epsilon,c+\epsilon,\ldots,c+\epsilon)^{\top}$. Clearly, we have that $d_{H}(X,Y)=\max_{i\in[3]}\min_{j\in[2]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\|\mathbf{v}_{3}-\mathbf{u}_{2}\|=\sqrt{d}\epsilon$. We also have that:

### v4

# arXiv:2505.24403v4[cs.LG] 27 Feb 2026

@@ 1 Introduction
Prior work estimates mainly the Lipschitz constant of architectures composed of fully-connected and convolutional layers. However, in several domains, input data might correspond to complex objects which consist of other simpler objects. We typically model these complex objects as sets or multisets (i. e., a generalization of a set). For instance, in computer vision, a point cloud is a set of data points in the 33-dimensional space. Likewise, in natural language processing, documents may be represented by multisets of word embeddings. Neural networks for sets typically consist of a series of fully-connected layers followed by an aggregation function which produces a representation for the entire multiset (Zaheer et al., 2017; Qi et al., 2017a). For the model to be invariant to permutations of the multiset’s elements, a permutation invariant aggregation function needs to be employed. Common functions include the sum, mean and max operators. While previous work has studied the expressive power of neural networks that employ the aforementioned aggregation functions (Wagstaff et al., 2019; 2022), their Lipschitz continuity and stability to perturbations remains underexplored. Besides those standard aggregation functions, other methods for embedding multisets have been proposed recently such as the Fourier Sliced-Wasserstein embedding which is bi-Lipschitz with respect to the Wasserstein distance (Amir and Dym, 2025).
@@ 3.1 Lipschitz Continuity of Aggregation Functions
The above theoretical result suggests that there is some correspondence between the three aggregation functions and the three distance functions for unordered multisets. In fact, each aggregation function seems to be closely related to a single metric. We also observe that while the Lipschitz constants of the sum and mean functions with respect to the matching distance and to EMD, respectively, is constant (equal to 11), the Lipschitz constant of the max function with respect to the Hausdorff distance depends on the dimension $d$ of the vectors (which is typically larger than 11). A direct consequence of the above Theorem is that no two of the considered distance functions are bi-Lipschitz equivalent.
@@ 4 Numerical Experiments
We experiment with two datasets from different domains: (i) ModelNet40: it contains 12,31112{,}311 3D CAD models that belong to $40$ object categories (Wu et al., 2015); and (ii) Polarity: it contains 10,66210{,}662 positive and negative labeled movie review snippets from Rotten Tomatoes (Pang and Lee, 2004). Note that the samples of both datasets can be thought of as multisets of vectors. Each sample of ModelNet40 is a multiset of 33-dimensional vectors. Polarity consists of textual documents, and each document is represented as a multiset of word vectors. The word vectors are obtained from a publicly available pre-trained model (Mikolov et al., 2013).
@@ 4.1 Lipschitz Constant of Aggregation Functions
In the first set of experiments, we empirically validate the results of Theorem 3.1 and Lemma 3.2. To obtain a collection of multisets of vectors, we train three different neural network models on the ModelNet40 and Polarity datasets. The difference between the three models lies in the employed aggregation function: mean, sum or max. More details about the different layers of those models are given in Appendix D. Note that the multisets used to verify the Lipschitz constants of the aggregation functions could, in principle, be generated by any means. We use those models to create multisets since the objective is to investigate how these functions behave in comparison to the derived bounds when the inputs are sampled from real distributions, rather than artificially generated data that do not occur in practice. Once the models are trained, we feed the test samples to them. For each test sample, we store the multiset of vectors produced by the layer of the model that precedes the aggregation function, and we also store the output of the aggregation function (a vector for each multiset). We then randomly choose $100$ test samples, and for each pair of those samples, we compute the EMD, Hausdorff distance and matching distance of their multisets of vectors, and also the Euclidean distance of their vector representations produced by the aggregation function. This gives rise to 99 combinations of distance functions and aggregation functions in total. Due to limited available space, we only show results for ModelNet40 in Figure 1. The results for Polarity can be found in Appendix E.1. Note that there are $\binom{100}{2}=4{,}950$ distinct pairs in total. Therefore, 4,9504{,}950 dots are visualized in each subfigure. To quantify the relationship between the output of the distance functions for multisets and the Euclidean distances of their vector representations, we compute and report the Pearson correlation coefficient. We observe from Figure 1 that the Lipschitz bounds (dash lines) successfully upper bound the Euclidean distance of the outputs of the aggregation functions. Note that all point clouds contained in the ModelNet40 dataset have equal cardinalities. Therefore, the conclusions of both Theorem 3.1 and Lemma 3.2 apply to this case, and thus we can derive Lipschitz constants for 77 out of the 99 combinations of distance functions for multisets and aggregation functions. We can see that the bounds that are associated with the mean and sum functions are tight, while those associated with the max function are relatively loose. We also observe that the distances of the representations produced by the mean and sum functions are very correlated with the distances produced by all three considered distance functions for multisets. On the other hand, the max function gives rise to representations that are less correlated with the produced distances.
@@ 4.2 Upper Bounds of Lipschitz Constants of Neural Networks for Sets
In the second set of experiments, we empirically validate the results of Theorem 3.4 and Lemma 3.5 on the ModelNet40 and Polarity datasets. We build neural networks that consist of three layers: (i) a fully-connected layer; (ii) an aggregation function; and (iii) a second fully-connected layer. Therefore, those models first transform the elements of the input multisets using an affine function, then aggregate the representations of the elements of each multiset and finally they transform the aggregated representations using a another affine function. Note that the Lipschitz constant of an affine function is equal to the largest singular value of the associated weight matrix, and can be exactly computed in polynomial time. We thus denote by $\text{Lip}(\text{FC}_{1})$ and $\text{Lip}(\text{FC}_{2})$ the Lipschitz constants of the two fully-connected layers, respectively. Note also that the Lipschitz constant of most activation functions (e. g., ReLU, LeakyReLU, Tanh) is equal to 11. Therefore, we can compute an upper bound of the Lipschitz constant of some models using Theorem 3.4 and Lemma 3.5. To train the models, we add a final layer to them which transforms the vector representations of the multisets into class probabilities. We use the same experimental protocol as in subsection 4.1 above (i. e., we randomly choose $100$ test samples). We only provide results for ModelNet40 in Figure 2, while the results for Polarity can be found in Appendix E.2. Since all point clouds contained in the ModelNet40 dataset have equal cardinalities, the conclusions of both Theorem 3.4 and Lemma 3.5 apply to this setting, and thus, once again, we can derive upper bounds on the Lipschitz constants for 77 out of the 99 combinations of distance functions for multisets and aggregation functions. We observe that the dash lines (Lipschitz upper bounds from Theorem 3.4 and Lemma 3.5) indeed upper bound the Euclidean distance of the outputs of the aggregation functions. We can also see that the bounds that are associated with the mean function are tight, while those associated with the sum and max functions are relatively loose and very loose, respectively. We also observe that the distances of the representations produced by the mean and sum functions are very correlated with the distances produced by EMD and matching distance. On the other hand, the max function gives rise to representations whose distances are less correlated with the distances produced by the distance functions for multisets.
@@ Appendix A Related Work
The seminal work of Zaheer et al. (2017) introduced DeepSets, a model that uses the sum aggregator to produce permutation-invariant set representations. They showed that when the elements of the input sets come from a countable set $\mathcal{X}$, then for an appropriate $f\colon\mathcal{X}\rightarrow\mathbb{R}$, the function defined as $g(\{x_{1},\ldots,x_{n}\})=\sum_{i=1}^{n}f(x_{i})$ maps the input sets injectively to $\mathbb{R}$. This result was later extended to multisets (Xu et al., 2019). In the countable case, an embedding dimension of 11 already suffices for injectivity. When $\mathcal{X}=\mathbb{R}$, multiset cardinalities are bounded by $m$ and $f$ is continuous, an embedding dimension of at least $m$ is both necessary and sufficient for injectivity (Wagstaff et al., 2019; 2022). For $\mathcal{X}=\mathbb{R}^{d}$, an embedding dimension of at least mdmd is necessary (Joshi et al., 2023). Most of these results rely on polynomial constructions to build injective multiset functions, after which the universal approximation theorem is invoked to argue that MLPs can approximate such polynomials. Amir et al. (2023) investigate whether MLP-based multiset functions are actually injective. They show that injectivity depends on the activation function. Specifically, analytic non-polynomial activation functions yield injective models, while networks with piecewise linear activation functions are injective only when $\mathcal{X}$ is finite or corresponds to certain irregular, countably infinite sets.
@@
Besides DeepSets, several other architectures and aggregation functions have been proposed recently. PointNet is another important architecture, primarily designed for point cloud data (Qi et al., 2017a). It consists of the same components as DeepSets, but instead of a sum aggregation function, it employs a max aggregator. To allow PointNet to capture local structures at different scales, Qi et al. (2017b) proposed PointNet++, a hierarchical model which applies PointNet recursively to nested partitions of the input set. Janossy pooling applies a neural network to all permutations of the input data and averages their outputs (Murphy et al., 2019). Since computing all permutations is generally intractable, the authors also propose some practical approximations. Set Transformer is a variant of the Transformer architecture designed for sets (Lee et al., 2019). Due to its attention mechanism, the model can capture interactions between elements in the input set. RepSet is another model designed for set-structured data which generates set representations by comparing the input set against learnable latent sets using a network flow algorithm (Skianis et al., 2020). FSPool sorts each feature across the elements of the set, and then computes a weighted sum of the elements where different weights can be learned for each feature dimension (Zhang et al., 2020). Pellegrini et al. (2021) propose a a learnable aggregation function which can approximate common aggregation functions (e. g., mean, sum, max), but also more complex functions. Kimura et al. (2024) introduce the Hölder’s Power DeepSets, a model tha generalized DeepSets by employing function known as power‑mean (a.k.a. Hölder mean), controlled by an exponent pp.
@@ Appendix B Proofs
Thus, we only need to prove that the triangle inequality holds. Let $Z\in\mathcal{S}(\mathbb{R}^{d}\setminus\{\mathbf{0}\})$. The three multisets can have different cardinalities. Let $|X|=m$, $|Y|=n$ and $|Z|=k$. Then, $X=\{\!\!\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{m}\}\!\!\}$, $Y=\{\!\!\{\mathbf{u}_{1},\mathbf{u}_{2},\ldots,\mathbf{u}_{n}\}\!\!\}$ and $Z=\{\!\!\{\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{k}\}\!\!\}$. There are 66 different cases. But it suffices to show that the triangle inequality holds when $|X|\geq|Y|\geq|Z|$, when $|Z|\geq|Y|\geq|X|$ and when $|X|\geq|Z|\geq|Y|$. Proofs for the rest of the cases are similar.
@@
Note that $d_{M}(X,Z)+d_{M}(Z,Y)$ can be written as a sum of $k$ terms, where each term corresponds to one of the above 44 sums of norms. If we take pairs of terms of types 22 and 33 and we sum them, we have that:
@@
Note also that type 22 occurs mnm-n times more than type 33. Therefore, using the inequalities for the 44 types of sums of norms above, we have:
@@ B.3.1 The mean function is Lipschitz continuous with respect to EMD
The mean function is thus Lipschitz continuous with respect to EMD and the Lipschitz constant is equal to 11.
@@ B.3.2 The mean function is not Lipschitz continuous with respect to the matching distance
Suppose that the mean function is Lipschitz continuous with respect to the matching distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>(2L+1)\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1}\}$ be two multisets, consisting of 22 and 11 vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$, and $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$. Clearly, we have that $d_{M}(X,Y)=\|\mathbf{v}_{2}\|=\sqrt{d}\epsilon$. We also have that:
@@ B.3.3 The mean function is not Lipschitz continuous with respect to the Hausdorff distance
Suppose that the mean function is Lipschitz continuous with respect to the Hausdorff distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>3L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ be two multisets, consisting of 33 and 22 vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(-c,-c,\ldots,-c)^{\top}$, $\mathbf{v}_{2}=\mathbf{u}_{2}=(c,c,\ldots,c)^{\top}$ and $\mathbf{v}_{3}=(c+\epsilon,c+\epsilon,\ldots,c+\epsilon)^{\top}$. Clearly, we have that $d_{H}(X,Y)=\max_{i\in[3]}\min_{j\in[2]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\|\mathbf{v}_{3}-\mathbf{u}_{2}\|=\sqrt{d}\epsilon$. We also have that: